# Causality Is Key to Understand and Balance Multiple Goals in Trustworthy ML and Foundation Models

### v4

@@ Abstract
Ensuring trustworthiness in machine learning (ML) systems is crucial as they become increasingly embedded in high-stakes domains. This paper advocates for integrating causal methods into machine learning to navigate the trade-offs among key principles of trustworthy ML, including fairness, privacy, robustness, accuracy, and explainability. While these objectives should ideally be satisfied simultaneously, they are often addressed in isolation, leading to conflicts and suboptimal solutions. Drawing on existing applications of causality in ML that successfully align goals such as fairness and accuracy or privacy and robustness, this paper argues that a causal approach is essential for balancing multiple competing objectives in both trustworthy ML and foundation models. Beyond highlighting these trade-offs, we examine how causality can be practically integrated into ML and foundation models, offering solutions to enhance their reliability and interpretability. Finally, we discuss the challenges, limitations, and opportunities in adopting causal frameworks, paving the way for more accountable and ethically sound AI systems.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Causal Trustworthy ML Cycle: Causal ML can leverage existing knowledge and causal auditing to enhance different components of trustworthiness: explainability, fairness, privacy, and accuracy while simultaneously advancing understanding through causal discovery.
@@
In recent years, machine learning (ML) has made remarkable strides, driving breakthroughs in natural language processing (Achiam et al., 2023), computer vision (Brooks et al., 2024), and decision-making systems (Jia et al., 2024). These advancements have led to widespread adoption across diverse domains, including healthcare (Singhal et al., 2025), finance (Lee et al., 2024), education (Modi et al., 2024), and social media (Bashiri & Kowsari, 2024), where ML models now play a crucial role in diagnostics, algorithmic trading, personalized learning, and content recommendation.
@@
Given their soaring influence, ensuring ethical and trustworthy ML systems has become a global priority (Lewis et al., 2020). Many international regulations and frameworks (European Commission, 2021; OECD, 2019; Group of Twenty (2019), G20; Infocomm Media Development Authority, 2020) seek to establish guidelines for fairness, explainability, robustness, and privacy protection. For the scope of our paper, we are aware of different definitions of trustworthiness, but will focus on five core dimensions that are both widely recognized and directly relevant to causal reasoning: fairness, privacy, robustness, explainability, and accuracy. We will introduce these dimensions and highlight their trade-offs and intersections below.
Fairness. Fairness in ML refers to the principle that systems should make unbiased decisions that do not discriminate against individuals or groups based on sensitive attributes such as race, gender, or socioeconomic status. ML systems have been shown to rely heavily on biased data, amplifying existing biases and leading to unequal outcomes (COMPAS, 2020). These systems often exhibit reduced accuracy for minority or underrepresented groups, further exacerbating disparities (Buolamwini & Gebru, 2018). Given the speed and scale of ML-enabled decisions, ensuring fairness is essential to prevent perpetuating and exacerbating societal inequalities at an unprecedented scale.
Privacy. Privacy in ML emphasizes the protection of individuals’ sensitive and personal data. It has been shown that even after removing identifiers such as names, information can still leak, and individuals can be reidentified through indirect attributes and data triangulation(Sweeney, 2000; Narayanan & Shmatikov, 2008; Ohm, 2010; Dwork, 2006). Additionally, sensitive information can be reconstructed from gradients during model training if data is not handled privately (Zhu et al., 2019; Geiping et al., 2020; Aono et al., 2017; Fredrikson et al., 2015). Privacy is crucial for ensuring compliance with data protection laws and safeguarding human rights. It also fosters trust for individuals to be more willing to contribute their data for model training if their safety and privacy were ensured.
Robustness. Robustness refers to the system’s ability to perform reliably under varying conditions, including adversarial attacks, noisy inputs, or distributional shifts. For example, models often underperform when faced with distribution shifts, such as changes in data characteristics between training and deployment environments (Hendrycks & Dietterich, 2019; Recht et al., 2019; Ovadia et al., 2019). Additionally, human-undetectable noise added to images can cause models to make incorrect predictions, highlighting their vulnerability (Szegedy et al., 2014; Goodfellow et al., 2015). Robustness is critical to ensuring the safety and reliability of AI systems, particularly in high-stakes applications such as healthcare and autonomous driving.
Explainability. Explainability refers to the ability of AI systems to provide clear and understandable reasoning behind their decisions or predictions. Deep neural networks (DNNs), often referred to as “black boxes,” are inherently complex and difficult to interpret, making them hard to audit and assess for fairness or correctness (Lipton, 2018; Doshi-Velez & Kim, 2017; Rudin, 2019). Explainability is closely tied to accountability, as it enables stakeholders to evaluate and challenge AI outputs when necessary. Furthermore, regulations such as the GDPR emphasize the “right to explanation,” which requires that individuals be informed about and understand how automated decisions affecting them are made (EC, 2016).
Trade-offs and Intersections. The trustworthy ML landscape involves complex trade-offs and interdependencies between key objectives such as fairness, privacy, accuracy, robustness, and explainability. Improving one aspect often comes at the expense of another, such as the trade-off between privacy and accuracy in differential privacy, where noise added to protect data reduces model accuracy (Xu et al., 2017; Carvalho et al., 2023). Similarly, achieving fairness frequently requires sacrificing predictive performance or resolving conflicts between competing fairness notions, such as demographic parity and equalized odds (Friedler et al., 2021; Kim et al., 2020). Trade-offs also arise in explainability and accuracy, as complex models like DNNs excel in performance but lack interpretability. Meanwhile, the relationship between fairness and privacy is nuanced, with evidence showing they can either conflict, as noise may lead to disparate outcomes, or complement each other by reducing bias (Pujol et al., 2020; Dwork et al., 2011).
Causality. One of the most influential causal frameworks is Pearl’s structural causal models (SCMs), which provide a systematic approach to reasoning about causality and integrating it into machine learning (Pearl, 2009). This framework defines causality as the relationship between the variables where a change in one variable (the cause) directly leads to a change in another variable (the effect). It establishes a directional and often mechanistic link, distinguishing relationships arising from mere correlations.
A key component of Pearl’s framework is the use of directed acyclic graphs (DAGs) and do-calculus, which offer a structured representation of causal dependencies and a formal method for performing causal inference. A causal DAG, denoted as $\mathcal{G}=(\mathbf{V},\mathcal{E})$, consists of a set of nodes $\mathbf{V}$ representing random variables and directed edges $\mathcal{E}$ encoding causal relationships among the variables.
Unlike correlation-based approaches, causality provides a framework for disentangling the underlying mechanisms that drive observed phenomena, offering a deeper interpretation of data. Causal frameworks have been successfully applied to audit and mitigate fairness (Kim et al., 2021; Kilbertus et al., 2017; Loftus et al., 2018) and to improve robustness (Schölkopf et al., 2022). The research about the connection between causality and privacy is still very limited, but some emerging studies show potential for applications (Tschantz et al., 2020). Finally, explainability is one of the core features of causality and comes pre-packaged with the causal framework. Despite the promising applications of causality for individual requirements of trustworthy AI, the potential to use causality to reconcile individual requirements of trustworthy ML remains largely under-explored.
Position. Despite significant advancements in research on individual dimensions of trustworthy ML such as fairness, privacy, and explainability—there is a notable lack of efforts to integrate these dimensions into a cohesive and unified framework. Each ethical principle addresses distinct challenges, yet their interplay often involves intricate trade-offs, particularly concerning model performance metrics such as accuracy. For example, mitigating fairness-related biases may require adjustments that compromise predictive precision, while enhancing explainability can impose constraints on model complexity. We argue that systematically addressing these trade-offs is a critical step toward developing AI systems that are both ethically sound and operationally efficient. While causality has been applied to address individual challenges such as fairness or interpretability, its potential to address the intersection of these challenges has largely been overlooked (see LABEL:app:relatedwork for a detailed review). In this paper, we argue that integrating causality into ML and foundation models offers a way to balance multiple competing objectives of trustworthy AI.
The structure of our paper is as follows. Section 2 analyzes how causality can reconcile multiple dimensions of trustworthy ML and explores how it can be integrated. Section 3 discusses how foundation models amplify existing ML trade-offs and introduce new challenges, for which we argue that causality provides a principled approach to overcoming these issues, and propose strategies for integrating causal reasoning into foundation models at different development stages. Section 4 covers limitations in applying causality to ML and foundation models and proposes future research directions, and Section 5 includes alternative views. Finally, Section 6 suggests key steps for advancing causality in ML and foundation models.
@@
Figure 2: While trustworthy AI involves inherent trade-offs between its key components, causality can help mitigate these tensions and enhance synergies.
@@ 2 Causality for Trustworthy ML
Trustworthy ML involves inherent trade-offs between core objectives such as accuracy, fairness, robustness, privacy, and explainability. Inevitable trade-offs can exist between accuracy and other objectives, fairness and privacy, and conflicting fairness notions. However, some other goals may reinforce each other, such as explainability aiding fairness assessment, and privacy enhancing robustness (Dwork & Lei, 2009; Hopkins et al., 2022).
Causality provides a principled approach to navigating these trade-offs by explicitly modeling data-generating processes and clarifying assumptions. This section first explores causal formulations for these trade-offs, and then introduce how causality can mitigate these tensions and support a more balanced approach to trustworthy ML.
@@ 2.1 Causality for Trade-offs in Trustworthy ML
Privacy vs. Accuracy. The differential privacy approach relies on adding noise to the data which is controlled by the parameter $\epsilon$ (the smaller value of $\epsilon$ corresponds to more noise, while the larger value indicates less noise and less privacy). Naturally, it hurts the accuracy of an algorithm learned on the privatized data. It is yet unknown how to avoid this fundamental trade-off between data protection and the utility of the data (Xu et al., 2017; Carvalho et al., 2023). One of ways how causality can inform privacy is provided by (Tschantz et al., 2020). The authors define privacy violations as causal effects, emphasizing that private information is leaked when an adversary can infer sensitive attributes from observable data due to causal pathways. Therefore, causal models can help identify, quantify, and mitigate such risks, offering a more systematic alternative to heuristic-based privacy measures.
By aligning privacy interventions with causal relationships, models can obscure sensitive attributes (e.g., sex, race) while preserving meaningful data dependencies, reducing the negative impact on accuracy. For example, causal graphs ensure that interdependent variables (e.g., age and education) are randomized together to avoid unrealistic combinations (e.g., “Age: 5; Education: Bachelor”). Preventing such inconsistencies not only improves accuracy but also reduces the likelihood of adversaries exploiting obfuscation patterns, enhancing overall privacy protection.
Fairness vs. Accuracy. Most of the statistical fairness literature focuses on improving fairness metrics while preserving accuracy as much as possible (Feldman et al., 2015; Wei & Niethammer, 2022; Wang et al., 2021). However, fairness often comes at the cost of reduced accuracy, as mitigating bias may require either obscuring predictive features that also contribute to discrimination or constraining model predictions within fairness-imposed boundaries (Pinzón et al., 2022; Cooper et al., 2021; Zliobaite, 2015; Zhao & Gordon, 2022).
A key issue is that many fairness-accuracy trade-offs arise from addressing correlations rather than causal relationships. Causal models can resolve these tensions by distinguishing legitimate predictive factors from spurious discriminatory pathways. By disentangling the direct and indirect effects of sensitive attributes on outcomes, causal interventions can mitigate unfair biases without sacrificing accuracy. For instance, counterfactual fairness ensures that individuals receive the same prediction regardless of their sensitive attributes in a counterfactual world where those attributes are altered (Kusner et al., 2017).
A compelling example comes from the COMPAS dataset, where Black defendants were more likely to be classified as high-risk for recidivism. Traditional statistical debiasing approaches treat race as a direct cause of the risk score, but a causal analysis reveals that increased recidivism risk is confounded by heightened policing in predominantly Black neighborhoods. By explicitly modeling this causal structure, fairness-enhancing interventions can adjust for the effect of over-policing, ensuring that predictions reflect true recidivism risk rather than biased enforcement patterns. This results in a more accurate and fairer risk assessment (Chiappa, 2019; Zafar et al., 2017).
@@
Conflicting Notions of Fairness. Fairness in ML is often constrained by conflicting definitions and measurement approaches. Friedler et al. (2021) highlight the fundamental tension between the “what you see is what you get” and “we are all equal” worldviews—where the former accepts disparities based on observed merit, while the latter seeks to correct historical inequalities. Causal graphs can crisply formulate different notions of fairness (Nabi & Shpitser, 2018; Chen et al., 2024), thus enabling feasible mitigation via path-specific causal effects (Avin et al., 2005).
@@
Kim et al. (2020) formalize fairness conflicts using the fairness-confusion tensor, showing that notions like demographic parity and equalized odds impose incompatible constraints. The causal approach mitigates these conflicts by focusing on fairness as a property of causal pathways rather than statistical dependencies (Rahmattalabi & Xiang, 2022). This allows for greater flexibility in aligning fairness interventions with real-world causal mechanisms, allowing better-informed choice of fairness metric.
Robustness vs. Accuracy. The trade-off between generalizability and accuracy is rooted in the observation that models trained to achieve high accuracy on a specific dataset often overfit to the peculiarities of that distribution. This overfitting compromises their ability to generalize to new, unseen distributions (Schölkopf et al., 2022). On the contrary, causal models focus on invariant relationships that hold across different environments, making them robust to distribution shifts. This robustness enhances the model’s ability to generalize to unseen data, improving accuracy in diverse settings. For example, causal representation learning disentangles stable causal factors, allowing the model to maintain performance when data distributions change. Moreover, Richens & Everitt (2024) prove that robust agents implicitly learn causal world models, further emphasizing the intrinsic interdependency between robustness and causality.
Explainability vs. Accuracy. Many complex algorithms, such as deep neural networks (DNN) or random forest (RF), have impressive predictive power but provide “black-box” solutions that are hard to question or evaluate (London, 2019; van der Veer et al., 2021). Causal models offer inherently interpretable structures by quantifying the contribution of each input feature to the output, providing clear, human-understandable explanations. Causal recourse further enhances explainabilit by offering actionable recommendations for individuals affected by model decisions, helping them achieve a more favorable outcome (Karimi et al., 2021).
@@
Fairness vs. Explainability. A particularly powerful approach within causal explainability is counterfactual explanations, which help users understand model decisions by asking “what if” questions. Counterfactual methods generate alternative scenarios where certain features are changed while keeping others constant, allowing for a direct assessment of how specific inputs influence predictions (Wachter et al., 2017; Karimi et al., 2020). Counterfactual explanations are particularly useful for fairness auditing as they can help identify why certain groups are adversely affected and guide corrective measures.
@@
In contrast, causal models inherently emphasize invariant relationships—patterns that are stable across various data distributions. Noise that disrupts non-causal relationships or spurious correlations can further enhance the robustness of these models to shifts in data. Finally, some results show, that causal models provide stronger guarantees for adversarial robustness with lower epsilon in differential privacy, thus allowing for lesser negative impact on accuracy (Tople et al., 2020).
Privacy vs. Fairness. Privacy mechanisms, such as noise addition, can disproportionately impact minority groups, leading to fairness concerns. Differentially Private Stochastic Gradient Descent (DP-SGD), for example, has been shown to degrade model accuracy more severely for underrepresented groups, exacerbating fairness disparities (Bagdasaryan et al., 2019). However, Causal models can guide privacy interventions by ensuring that noise is applied in ways that do not disrupt fairness-critical relationships. For instance, a causal graph can reveal which features or pathways should be preserved to maintain fairness while protecting privacy.
Prediction Accuracy vs. Intervention Accuracy. One of the key advantages of the causal framework is its ability to support not just prediction but also intervention (Hernán & Robins, 2020; Schölkopf et al., 2022). While predictive models are sufficient in some domains, many high-stakes applications—such as healthcare, policy-making, and personalized treatment—require actionable interventions. In these settings, understanding causal relationships is essential, as the objective is not only to predict outcomes but also to influence them.

### v5

@@ Abstract
Ensuring trustworthiness in machine learning (ML) systems is crucial as they become increasingly embedded in high-stakes domains. This paper advocates for integrating causal methods into machine learning to navigate the trade-offs among key principles of trustworthy ML, including fairness, privacy, robustness, accuracy, and explainability. While these objectives should ideally be satisfied simultaneously, they are often addressed in isolation, leading to conflicts and suboptimal solutions. Drawing on existing applications of causality in ML that successfully align goals such as fairness and accuracy or privacy and robustness, this position paper argues that the causal approach is essential for balancing multiple competing objectives in both trustworthy ML and foundation models. Beyond highlighting the potential to resolve trade-offs in trustworhy ML, we examine how causality can be practically integrated into ML and foundation models. Finally, we discuss the challenges, limitations, and opportunities in adopting causal frameworks, paving the way for more accountable and ethically sound ML systems.
@@ 1 Introduction
In recent years, machine learning (ML) has made remarkable strides, driving breakthroughs in natural language processing, computer vision, and decision-making systems jia2024decision ; achiam2023gpt . These advancements have led to widespread adoption across diverse domains, including healthcare singhal2025toward , finance lee2024survey , education team2024learnlm , and social media bashiri2024transformative , where ML models now play a crucial role in diagnostics, algorithmic trading, personalized learning, and content recommendation.
Given their soaring influence, ensuring ethical and trustworthy ML systems has become a global priority lewis2020global . Many international regulations and frameworks EU_AI_Act_2021 ; OECD_AI_Principles_2019 ; G20_AI_Principles_2019 ; Singapore_AI_Governance_2020 seek to establish guidelines for fairness, explainability, robustness, and privacy protection.
@@
Various organizations provide slightly different definitions and components of trustworthy ML. For the scope of our paper, we focus on five core dimensions that are both widely recognized and directly relevant to causal reasoning: fairness, privacy, robustness, explainability, and accuracy. We will introduce these dimensions and highlight their trade-offs and intersections below.
@@
Figure 1: While trustworthy ML involves inherent trade-offs between its key components, causality can help mitigate these tensions and enhance synergies.
Trustworthy ML must balance competing objectives like accuracy, fairness, robustness, privacy, and explainability. While trade-offs—such as between accuracy and fairness or privacy—are often unavoidable, some goals can be mutually reinforcing (e.g., explainability aiding fairness, privacy supporting robustness) dwork2009differential ; hopkins2022robustness .
Fairness. Fairness in ML refers to the principle that systems should make unbiased decisions that do not discriminate against individuals or groups based on sensitive attributes such as race, gender, or socioeconomic status. ML systems have been shown to rely heavily on biased data, amplifying existing biases and leading to unequal outcomes COMPAS . These systems often exhibit reduced accuracy for minority or underrepresented groups, further exacerbating disparities buolamwini2018gender . Given the speed and scale of ML-enabled decisions, ensuring fairness is essential to prevent perpetuating and exacerbating societal inequalities at an unprecedented scale.
Privacy. Privacy in ML emphasizes the protection of individuals’ sensitive and personal data. It has been shown that even after removing identifiers such as names, information can still leak, and individuals can be reidentified through indirect attributes and data triangulationsweeney2000simple ; narayanan2008robust ; ohm2010broken ; dwork2006differential . Additionally, sensitive information can be reconstructed from gradients during model training if data is not handled privately zhu2019deep ; geiping2020inverting ; aono2017privacy ; fredrikson2015model . Privacy is crucial for ensuring compliance with data protection laws and safeguarding human rights. It also fosters trust for individuals to be more willing to contribute their data for model training if their safety and privacy are ensured.
Robustness. Robustness refers to the system’s ability to perform reliably under varying conditions, including adversarial attacks, noisy inputs, or distributional shifts. For example, models often underperform when faced with distribution shifts, such as changes in data characteristics between training and deployment environments hendrycks2019benchmarking ; recht2019imagenet ; ovadia2019can . Additionally, human-undetectable noise added to images can cause models to make incorrect predictions, highlighting their vulnerability szegedy2014intriguing ; goodfellow2015explaining . Robustness is critical to ensuring the safety and reliability of AI systems, particularly in high-stakes applications such as healthcare and autonomous driving.
Explainability. Explainability refers to the ability of AI systems to provide clear and understandable reasoning behind their decisions or predictions. Deep neural networks (DNNs), often referred to as “black boxes,” are inherently complex and difficult to interpret, making them hard to audit and assess for fairness or correctness lipton2018mythos ; doshi2017towards ; rudin2019stop . Explainability is closely tied to accountability, as it enables stakeholders to evaluate and challenge AI outputs when necessary. Furthermore, regulations such as the GDPR emphasize the “right to explanation,” which requires that individuals be informed about and understand how automated decisions affecting them are made GDPR_site .
Trade-offs and Intersections. The trustworthy ML landscape involves complex trade-offs and interdependencies between key objectives such as fairness, privacy, accuracy, robustness, and explainability ( Figure 1). Improving one aspect often comes at the expense of another, such as the trade-off between privacy and accuracy in differential privacy, where noise added to protect data reduces model accuracy xu2017privacy ; carvalho2023towards . Similarly, achieving fairness frequently requires sacrificing predictive performance or resolving conflicts between competing fairness notions, such as demographic parity and equalized odds friedler2021possibility ; kim2020fact . Trade-offs also arise in explainability and accuracy, as complex models like DNNs excel in performance but lack interpretability. Meanwhile, the relationship between fairness and privacy is nuanced, with evidence showing they can either conflict, as noise may lead to disparate outcomes, or complement each other by reducing bias pujol2020fair ; dwork2011fairness .
Causality. One of the most influential causal frameworks is Pearl’s structural causal models (SCMs), which provide a systematic approach to reasoning about causality and integrating it into machine learning pearl_causality_2009 . This framework defines causality as the relationship between the variables where a change in one variable (the cause) directly leads to a change in another variable (the effect). It establishes a directional and often mechanistic link, distinguishing relationships arising from mere correlations.
A key component of Pearl’s framework is the use of directed acyclic graphs (DAGs) and do-calculus, which offer a structured representation of causal dependencies and a formal method for performing causal inference. While important for causal reasoning, recent advances have proposed approximate methods for applying causality with partial causal graphs or discover implicit causal graphs from the data, making causal approaches more applicable in practice kaddour2022causal ; burgess2018understanding ; kim2021counterfactual ; russell2017worlds ; ehyaeibridging ; zuo2022counterfactual .
Unlike correlation-based approaches, causality provides a framework for disentangling the underlying mechanisms that drive observed phenomena, offering a deeper interpretation of data. Causal frameworks have been successfully applied to audit and mitigate fairness kim2021counterfactual ; kilbertus2017avoiding ; loftus18 and to improve robustness scholkopf2022causality . The research about the connection between causality and privacy is still very limited, but some emerging studies show potential for applications tschantz2020sok . Finally, explainability is one of the core features of causality and comes pre-packaged with the causal framework. Despite the promising applications of causality for individual requirements of trustworthy AI, the potential to use causality to reconcile multiple requirements of trustworthy ML remains largely under-explored.
Position. Despite significant advancements in research on individual dimensions of trustworthy ML such as fairness, privacy, and explainability—there is a notable lack of efforts to integrate these dimensions into a cohesive and unified framework. Each ethical principle addresses distinct challenges, yet their interplay often involves intricate trade-offs, particularly concerning model performance metrics such as accuracy. For example, mitigating fairness-related biases may require adjustments that compromise predictive precision, while enhancing explainability can impose constraints on model complexity. We argue that systematically addressing these trade-offs is a critical step toward developing AI systems that are both ethically sound and operationally efficient. While causality has been applied to address individual challenges such as fairness or interpretability, its potential to address the intersection of these challenges has largely been overlooked. In this position paper, we argue that integrating causality into ML and foundation models offers a way to balance multiple competing objectives of trustworthy AI. We base our position on a review of how causal machine learning has advanced various aspects of trustworthy ML. From this, we identify key mechanisms behind objective trade-offs and show how causal methods can help resolve them. We then explore efforts to integrate causality into foundation models, highlighting unique challenges and outlining future directions for applying causal reasoning in this new context.
The structure of our paper is as follows. Section 2 discusses how causality can reconcile multiple dimensions of trustworthy ML and explores how it can be integrated. Section 3 discusses how FMs amplify existing ML trade-offs and introduce new challenges and propose causal approach to overcoming these issues as well as strategies for integrating causal reasoning into FMs. Section 4 covers limitations and possible solutions, and Section 5 includes alternative views. Finally, Section 6 suggests actionable steps for advancing causality in ML and FMs and proposes future research directions.
@@ 2 Causality for Trustworthy ML
Causal modeling provides a principled way to address these tensions by capturing the underlying data-generating processes. This section first explores how causal approaches can help reconcile key trade-offs, then outlines methods for integrating causality into machine learning.
@@ 2.1 Causality for Trade-offs in Trustworthy ML
Fairness vs. Accuracy. Efforts to improve fairness in ML often reduce accuracy, as they may obscure predictive features or constrain outputs to meet fairness criteria pinzon2022impossibility ; cooper2021emergent ; zliobaite2015relation ; zhao2022inherent . Many such trade-offs stem from focusing on correlations rather than underlying causes. Causal models offer a principled solution by isolating spurious pathways from valid predictive factors. For example, counterfactual fairness ensures predictions remain stable when sensitive attributes (e.g., race) are altered in a counterfactual world kusner2017counterfactual . A compelling example comes from the COMPAS dataset, where Black defendants were more likely to be classified as high-risk for recidivism. While statistical methods attribute this to race, a causal model might reveal over-policing in Black neighborhoods as a confounder of recidivism predictions neil2023racial . Adjusting for such causal pathways allows for risk assessments that are both more accurate and fair chiappa2019path ; zafar2017fairness . Causal frameworks thus mitigate bias without discarding valid predictive signals, effectively softening the fairness-accuracy trade-off.
@@
Conflicting Notions of Fairness. Fairness in ML is often constrained by conflicting definitions and measurement approaches. friedler2021possibility highlight the fundamental tension between the “what you see is what you get" and “we are all equal" worldviews—where the former accepts disparities based on observed merit, while the latter seeks to correct historical inequalities. Causal graphs can crisply formulate different notions of fairness nabi2018fair ; 3a2e333502f5035f0f78318c7722ec1a7b7265bb , thus enabling feasible mitigation via path-specific causal effects avin2005identifiability .
@@
kim2020fact formalize fairness conflicts using the fairness-confusion tensor, showing that notions like demographic parity and equalized odds impose incompatible constraints. The causal approach mitigates these conflicts by focusing on fairness as a property of causal pathways rather than statistical dependencies rahmattalabi2022promises . This allows for greater flexibility in aligning fairness interventions with real-world causal mechanisms, allowing better-informed choice of fairness metric.
@@
Explainability for Fairness. A particularly powerful approach within causal explainability is counterfactual explanations, which help users understand model decisions by asking “what if" questions. Counterfactual methods generate alternative scenarios where certain features are changed while keeping others constant, allowing for a direct assessment of how specific inputs influence predictions wachter2017counterfactual ; karimi2020model . Counterfactual explanations are particularly useful for fairness auditing as they can help identify why certain groups are adversely affected and guide corrective measures.
Privacy vs. Accuracy. Differential privacy introduces noise—controlled by parameter $\epsilon$—to protect sensitive data, but this often reduces model accuracy xu2017privacy ; carvalho2023towards . The trade-off between privacy and utility remains fundamentally unresolved. Causal models offer a more principled alternative by framing privacy breaches as causal effects—where adversaries infer sensitive attributes through observable pathways tschantz2020sok .
By aligning interventions with causal structure, models can obscure private information (e.g., race, sex) while preserving essential dependencies, minimizing accuracy loss. For instance, causal graphs can enforce realistic variable combinations (e.g., aligning age and education), preventing implausible data patterns and reducing the risk of leakage. This approach improves both privacy protection and model fidelity.
Privacy vs. Fairness. Privacy mechanisms, such as noise addition, can disproportionately impact minority groups, leading to fairness concerns. Differentially Private Stochastic Gradient Descent (DP-SGD), for example, has been shown to degrade model accuracy more severely for underrepresented groups, exacerbating fairness disparities bagdasaryan2019differential . However, Causal models can guide privacy interventions by ensuring that noise is applied in ways that do not disrupt fairness-critical relationships. For instance, a causal graph can reveal which features or pathways should be preserved to maintain fairness while protecting privacy.
@@
In contrast, causal models inherently emphasize invariant relationships—patterns that are stable across various data distributions. Noise that disrupts non-causal relationships or spurious correlations can further enhance the robustness of these models to shifts in data. Finally, some results show, that causal models provide stronger guarantees for adversarial robustness with lower epsilon in differential privacy, thus allowing for lesser negative impact on accuracy tople2020alleviating .
Robustness vs. Accuracy. The trade-off between generalizability and accuracy is rooted in the observation that models trained to achieve high accuracy on a specific dataset often overfit to the peculiarities of that distribution. This overfitting compromises their ability to generalize to new, unseen distributions scholkopf2022causality . On the contrary, causal models focus on invariant relationships that hold across different environments, making them robust to distribution shifts. This robustness enhances the model’s ability to generalize to unseen data, improving accuracy in diverse settings. For example, causal representation learning disentangles stable causal factors, allowing the model to maintain performance when data distributions change. Moreover, richens2024 prove that robust agents implicitly learn causal world models, further emphasizing the intrinsic interdependency between robustness and causality.
Explainability vs. Accuracy. Many complex algorithms, such as deep neural networks (DNN) or random forest (RF), have impressive predictive power but provide “black-box" solutions that are hard to question or evaluate london2019artificial ; van2021trading . Causal models offer inherently interpretable structures by quantifying the contribution of each input feature to the output, providing clear, human-understandable explanations. Causal recourse further enhances explainability by offering actionable recommendations for individuals affected by model decisions, helping them achieve a more favorable outcome KarSchVal21 .
Prediction Accuracy vs. Intervention Accuracy. One of the key advantages of the causal framework is its ability to support not just prediction but also intervention hernan2020causal ; scholkopf2022causality . While predictive models are sufficient in some domains, many high-stakes applications—such as healthcare, policy-making, personalized treatment, as well as addressing structural biases and fairness, require actionable interventions kusner2018causal ; kusner2019making . In these settings, understanding causal relationships is essential, as the objective is not only to predict outcomes but also to influence them.